IhelpUuHELPme:

William is putting up signs on the edge of the road every 1/16 of a mile. He just put a sign down and he has 7/8 of a mile of road left in which to place his signs. He wants to determine how many more signs he should put on the edge of the road. William should put up ________ more signs.

MiraAngel:

@jhonyy9 May you try to help with this question when you have time?... I'm kind of brain dead at the moment.

IhelpUuHELPme:

oof

jhonyy9:

how many of 1/16 give to you 1 mile ? how many of 1/16 make 1 integer ? how many time you need assuming the 1/16 to result 1 ?

xxcoffee:

ye im dead rn yk my brain had to much ice cream and im finna have cake so yk

MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 how many of 1/16 give to you 1 mile ? how many of 1/16 make 1 integer ? how many time you need assuming the 1/16 to result 1 ? \(\color{#0cbb34}{\text{End of Quote}}\) Would all of them be 16?

jhonyy9:

yeah

xxcoffee:

i feel dumb i did da work and got 15 i did it wrong oop

jhonyy9:

bc. 16*(1/16) =16/16 = 1

MiraAngel:

Okay

MiraAngel:

Half of 1/16 is 7/8 right?

MiraAngel:

no... it's not. scratch that

MiraAngel:

14/16 is 7/8 doubled

jhonyy9:

1/16/2 = 1/16 *1/2 = 1/32

MiraAngel:

So two more signs to go?

IhelpUuHELPme:

so 16

IhelpUuHELPme:

nope it was 14

jhonyy9:



MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IhelpUuHELPme so 16 \(\color{#0cbb34}{\text{End of Quote}}\) No. Some signs have already been placed. 16 is the total of how many should be placed... We are trying to figure out how many more

IhelpUuHELPme:

i kept the answer in already

IhelpUuHELPme:

it was 14

IhelpUuHELPme:

i really need to be more patient

MiraAngel:

Okay, so two were already placed and 14 more to go. Let jhonny or me know if you have more questions.

IhelpUuHELPme:

thank you

IhelpUuHELPme:

how do i tag people

IhelpUuHELPme:

by the way

jhonyy9:

in this way @jhonyy9

IhelpUuHELPme:

i dont understand

jhonyy9:

just write this sign of @ and the nickname of who wan taging

IhelpUuHELPme:

cool

IhelpUuHELPme:

thank you

jhonyy9:

yw anytime

IhelpUuHELPme:

thank you jhonny and mira and kathy for trying to help me

IhelpUuHELPme:

i will tag yall if i need any more help

jhonyy9:

ok bye bye

IhelpUuHELPme:

bye

MiraAngel:

Okay! Thank you! Good luck!